How to Format the Paper for the ICNNAI-2010

Cyrano Daich

(Do not show your name when to be submitted! Should be appeared only for camera-ready)

Brest State Technical University Moskowskaja 267, Brest 224017 Republic of Belarus akira@bsty.by

(Do not show your affiliation either when to be submitted.)

Abstract

In this sample, we show how to format your paper to be submitted to the ICNNAI-2010. You may compile simply by replacing texts, figures, tables and so on in this sample.tex with your orignal ones, following the instractions below. Regarding Abstract, needless to be mentioned but, the following rule is requested. (i) it should be made up of only one paragraph; (ii) it should be independent from the text (e.g., avoid refering to the paper in the reference list, or, if you show an abbrebiation such as Genetic Algorithm (GA), you may repeat this in its first appearance in the text, and so on).

1 INTRODUCTION

First of all, please notice that all papers received will be reviewed following a blind reviewing policy (the reviewers will not know the identities of the authors). So, please make your submission anonymous by removing all references to the authors. That is, please do not include any information with which the reviewer could identify the authors – leave the authors and their affiliations line above empty, and obvious self-references must not be mentioned anywhere thereafter in the submission.

The title of the paper should be written as follows. All words in the title except for particles (i.e., “the” “an” and “a”) and prepositions (“of” “for”...) should be started with upper case letter. See the title of this example.

The title of all the sections should be all with upper case letters. Each section may be included subsections and sub-sub-sections if necessary. The title of sub-sections and sub-sub-sections, if any, should start with upper case letter following lower case letters. See the following example.

1.1 No indent policy

Start all sections, sub-sections, and sub-sub-sections without indent as already seen in this sample.

1.1.1 Blank line between two paragraphs

As already shown at the beginning of this section all paragraphs should be separated from the previous one with one empty line, instead of starting with indent.

1.2 Variables in text

Variables in text should be italisized like x, y, or z = f(t).

1.3 Equations

Equations should be centered as follows. Don’t forget to put “.” if the equation conclude a sentence. Or “,” if it’s appropriate.

s_{i}(t + 1) = sgn (∑_{j≠i}^{N} w_{ij}s_{j}(t)) .

Or, if necessary, you may number them like:

w_{ij}(t + 1) = w_{ij}(t) + ηx_{j}(t)y_{i}(t).  (1)

1.4 Citations

Citation should be (author, year) style. Or author (year) depending on the situation.

Basically citation can be categorized into (i) from books; (ii) from conference proceedings; (iii) from